OpenStudy (anonymous):

3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1)

OpenStudy (deadpool):

multiply each number outside the parenthesis into the parenthesis

OpenStudy (deadpool):

3(x-2) 3*x,3*-2 3x,-6 so its now 3x-6

OpenStudy (anonymous):

i know how to do it but its a fraction so it confuses me

OpenStudy (deadpool):

do all the work and then type it so i can see

OpenStudy (anonymous):

11x-30=22x-16

OpenStudy (deadpool):

ok now subtract11x and then add 16 11x-11x-30+16=22x-11x-16+16

OpenStudy (anonymous):

-14=11x

OpenStudy (deadpool):

ok now divide 11 from both sides

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

but thats nopt a fraction

OpenStudy (deadpool):

no no 14/11=x

OpenStudy (anonymous):

-14?

OpenStudy (deadpool):

thats what i meant but you are dividing 11 not subtracting